OpenStudy (anonymous):

write an equation of a circle with the center at the origin and a radius of 1/2

OpenStudy (anonymous):

\[x^{2}+y{^2}=1/4\]

OpenStudy (campbell_st):

the general form of the circle is (x -h)^2 + (y - k)^2 = r^2 (h, k) is the centre and r = radius in your question h = 0, k = 0 and r = 1/2

OpenStudy (anonymous):

thanks guys